# Limitations of Existing Work, and Guidelines for Future Research

### v1

@@ Abstract
Despite location being increasingly used in decision-making systems employed in many sensitive domains such as mortgages and insurance, astonishingly little attention has been paid to unfairness that may seep in due to the correlation of location with characteristics considered protected under anti-discrimination law, such as race or national origin. This position paper argues for the urgent need to consider fairness with respect to location, termed spatial fairness, by outlining the harms that continue to be perpetuated due to location’s correlation with protected characteristics. This interdisciplinary work connects knowledge from fields such as public policy, economic development, and geography to highlight how fair-AI research currently falls short of correcting for spatial biases, and does not consider challenges unique to spatial data. Furthermore, we identify limitations of the handful of spatial fairness work proposed so far, and finally, detail guidelines for future research so subsequent work may avoid such issues and help correct spatial biases.
fairness machine learning, fair-AI, machine learning, ICML
@@ 1 Introduction
The adoption of data-driven decision-making systems has skyrocketed across the board in the last two decades due to their ease of deployment and proficiency in learning patterns from enormous amounts of historical data to make decisions, reducing human effort while increasing efficiency. As the fairness in artificial intelligence (fair-AI) literature (Mehrabi et al., 2021; Pessach & Shmueli, 2022) has noted, however, such decision-making systems can also replicate undesirable trends in historical data, leading to undesirable patterns such as historic racial or gender discrimination being perpetuated further, or worse–possibly even being amplified.
@@
This observation has led to extensive research on developing fairer AI systems, which has centered on protecting against discrimination with respect to personal characteristics or traits protected by anti-discrimination law, commonly referred to in the literature as (legally) protected characteristics, traits, or attributes. For example, the Fair Housing Act of 1968 in the U.S. protects against discrimination based on race, religion, sex, disability, national origin, and familial status in the housing domain (Mehrabi et al., 2021; Lee & Floridi, 2021). There is similar legislation in the U.S. and other parts of the world (e.g., E.U.) protecting against discrimination on the basis of legally protected characteristics in multiple domains, particularly employment, credit, and housing (Xiang & Raji, 2019). However, this singular focus in the computer science community on legally protected characteristics ignores the possibility that other, seemingly neutral characteristics may be correlated to them. For example, research has documented discrimination against racial minorities in the mortgage industry (Pager & Shepherd, 2008; Ross et al., 2008; Massey et al., 2016), despite the Equal Credit Opportunity Act prohibiting creditors from discriminating against applicants on the basis of race (Mehrabi et al., 2021). This discrimination likely occurs not due to an intent to discriminate but due to the presence of the extensive correlation of race with zip codes in the United States due to prejudiced practices such as redlining prevalent in the past (Schill & Wachter, 1994; Woods, 2012).
As such examples show, simply not considering protected characteristics in decision-making, an approach known as ‘fairness through unawareness’ (Grgic-Hlaca et al., 2016; Kusner et al., 2017) is insufficient in practice (Lee & Floridi, 2021). Other fair-AI work has argued to “repair” or fix attributes that may act as proxies for legally protected characteristics (He et al., 2020; Kamiran et al., 2010; Berk et al., 2017; Zafar et al., 2017). However, subsequent research has shown this is ineffective in practice (Corbett-Davies & Goel, 2018; Lee & Floridi, 2021). Other work argues for dropping both the proxies and protected characteristics from the data altogether (Datta et al., 2017), but this is not ideal since a proxy attribute can carry information relevant to decision-making, such as crime rates in zip codes for mortgage decisions.
@@
While many attributes may be correlated with a legally protected attribute, the use of location is particularly troubling due to its potential to be correlated with multiple legally protected characteristics, such as race or national origin (Marshall, 1974; Willborn, 1984; Romei & Ruggieri, 2014; Becker, 2010). Table 1 outlines legally protected characteristics observed to have significant correlation with location, or can be inferred from location data. Extensive research in public policy, economic development, and sociology has documented the extensive spatial segregation that still adversely affects society today. The danger that such historical biases that predominantly disadvantage racial and ethnic minorities (Rothstein, 2015; Schill & Wachter, 1994; Rothstein, 2017) may seep into outcomes of decision-making systems is alarming and warrants greater attention.
Fairness concerning location is further complicated by challenges specific to spatial data, limiting the generalizability of current fair-AI methods, motivating the development of fairness definitions and methods tailored for spatial fairness.
@@
This work advocates for the need for a focus on spatial fairness within fair-AI. We use the interplay of legislation, historical discrimination, and AI systems within the U.S. to motivate greater attention towards bias due to location throughout the paper, however, the issue of discrimination due to location because of its correlation with other attributes is not limited to the U.S. alone (Galster & Friedrichs, 2015; Korsu & Wenglenski, 2010). By drawing on literature in the law, sociology, and economic development as well as computer science, we aim to ensure that future research by computer scientists in fair-AI does not fall into the trap of not being implementable in practice due to its noncompliance with legal or other standards (Xiang & Raji, 2019; Barocas & Selbst, 2016; Kim, 2022). We further delineate challenges specific to location that future work would need to be aware of, and finally draw on relevant research in fields such as public policy and the law to define guidelines for subsequent work in spatial fairness. Our main contributions are:
@@
  • Highlight the importance of systematically addressing spatial fairness within the context of fair-AI, drawing upon scholarly literature in the domains of law and public policy
@@
  • Draw from fields of public policy, law, transport geography and economic development to delineate guidelines for future work on spatial fairness
  • Motivate interdisciplinary work that comprehensively brings together relevant knowledge from different domains to push fair-AI forward and address a blind spot in current fair-AI research
The paper is organized as follows. Section 2 motivates spatial fairness by disentangling the complex background on spatial biases and its ongoing societal impact, using growing legal efforts to establish protection against discrimination based on zip codes as a case study. Challenges unique to spatial data that limit generalizability of current fair-AI work to spatial contexts are detailed in Section 3. Section 4 outlines limitations of the handful of existing spatial fairness definitions and techniques. Guidelines for future spatial fairness work to avoid such pitfalls and comply with legal and public policy standards are delineated in Section 5. We conclude with Section 6.
@@
[…]
Protected in housing in 15 states (pol, )
[…]
Table 1: Protected characteristics in housing (federal: Fair Housing Act; state: Policy Surveillance Program), credit (Equal Credit Opportunity Act, and employment (Equal Employment Opportunity Commission. (Mehrabi et al., 2021; Chen et al., 2019; pol, )
@@ 2 Context and Motivation: Location bias
We discuss the history of spatial segregation within the U.S., resulting in pervasive location-based bias and discrimination even today, and outline the case for legal protection for discrimination on the basis of location, particularly zipcodes.
@@ 2.1 Spatial unfairness in the real world
Although the fair-AI literature has yet to explore bias due to location in depth, it has been extensively covered in fields such as public policy, urban planning, and the law (Schill & Wachter, 1994; Rothstein, 2017; Baum et al., 2015). For example, neighborhoods in the United States have been historically correlated with race, a legally protected characteristic in many domains (e.g. housing, employment, credit), in complex ways due to prejudiced practices such as redlining [^1] Redlining is the discriminatory practice of systematically denying individuals access to mortgages, insurance, and other financial services living in certain neighborhoods, based on their race or ethnicity. (Woods, 2012; Rothstein, 2017). Even though the practice of redlining was prohibited by the Fair Housing Act in 1968 [42 U.S.C. §§ 3601 - 3619], unfortunately, its effects continue to linger (Angwin et al., 2017; Mehrabi et al., 2021). Zip codes in many parts of the country still correlate with race, potentially acting as a proxy for race in data-driven decision-making scenarios where zip codes are considered as a feature, such as in mortgages and insurance. This has been frequently documented. For example, the non-profit ProPublica showed that drivers who live in (racial) minority neighborhoods are charged higher insurance premiums than drivers of same risk levels who reside in predominantly white neighborhoods (Angwin et al., 2017). Research has also found that prices for ride-hailing trips in the city of Chicago from or to neighborhoods with larger non-white populations or higher poverty levels were significantly likely to be charged higher prices (Pandey & Caliskan, 2021; Saxena et al., 2023b).
@@
Neighborhoods (and their zip codes) may also be correlated with immigrants, i.e., neighborhoods where immigrants are more likely to live, which can act as a proxy for national origin, another characteristic protected against discrimination in domains such as housing, credit, and employment under U.S. law (Mehrabi et al., 2021). Furthermore, neighborhoods may also be correlated with income (Mode et al., 2016). In addition to the ride-hailing example mentioned above which found higher trip prices significantly associated with neighborhoods with higher poverty levels, residents of low-income neighborhoods may also be subjected to other forms of income-related discrimination. For example, families receiving housing vouchers often face barriers in moving to lower-poverty and higher-income neighborhoods due to discrimination by landlords against their source of income (Mazzara & Knudsen, 2019). While not illegal at the federal level, some states have laws protecting people from discrimination based on source of income (Table 1). As a result, many families receiving housing vouchers still end up in low-income, racially segregated neighborhoods despite being able to afford homes in better neighborhoods (Rothstein, 2015). Such families might suffer doubly if they are also then subjected to spatial discrimination due to their low-income, racially segregated neighborhood.
@@
Despite racial segregation in housing being banned in the U.S. in 1968, spatial segregation by race has persisted, resulting in “racially homogenous public institutions that are geographically defined,” (Fulwood III, 2017). As a consequence, such neighborhoods have reduced access to economic, educational, and social opportunities (Chetty et al., 2016; Li et al., 2013), which further hinders future prospects for residents in those neighborhoods. This also manifests in house prices, with a study finding that the racial composition of a neighborhood was a stronger determinant of a home’s appraised value in 2015 than in 1980 (Howell & Korver-Glenn, 2021). Furthermore, research has shown that prices for houses in majority-white neighborhoods are significantly higher than for similar houses in neighborhoods that happen to be majority-black, worsening the wealth gap between whites and racial minorities, especially African Americans, since a home is typically a family’s most significant financial asset according to a report by Pew Research Center (Taylor et al., 2011). There has been no significant change over the past few decades in the manner in which majority-black neighborhoods continue to be encircled by spatial disadvantage (Sharkey, 2014). Ethnographic research has documented that even middle class African-American families with an annual income of at least \$100,000 tend to live in neighborhoods with the same disadvantages as the average white family with an annual income of \$30,000 or lower (Sharkey, 2014).
@@
While we focus on anti-discrimination law in the United States to motivate this problem, unfortunately such spatial inequities are not restricted to the U.S. alone. Research in fields such as economics and public policy have documented rising spatial inequities in countries across the world (Kanbur & Venables, 2005; Kochendörfer-Lucius & Pleskovic, 2009; Zhao & Tong, 2000; Feldman & Storper, 2018).
@@ 2.2 Legal support for spatial fairness
The law in the United States recognizes two forms of discrimination for protected characteristics or classes (Lee & Floridi, 2021; Baum et al., 2015): disparate treatment and disparate impact. In layman terms, disparate treatment can be thought of as direct or intentional discrimination while disparate impact can be viewed as indirect or unintentional discrimination (Barocas & Selbst, 2016; Primus, 2010). While disparate treatment focuses on the intent to discriminate, disparate impact focuses on outcomes that are discriminatory towards members of a protected class despite a seemingly neutral decision-making policy.
@@
For both forms of discrimination, U.S. law provides protections against discrimination on the basis of legally protected characteristics or attributes in domains such as housing [42 U.S.C. § 3601 et seq], credit [15 U.S.C. 1691 et seq.], and employment [42 U.S.C. § 2000e]. Some states define additional protected attributes, for example, marital status is also a protected class in the state of California in the domain of employment (Gelb & Frankfurt, 1982) although it is a protected characteristic only in the domain of credit at the federal level. Typically under U.S. anti-discrimination law, protected characteristics are those which are considered immutable, i.e., characteristics that were not chosen and cannot be changed over time (Hoffman, 2010), such as race, color, and sex. Race and color were among the first classes to be considered immutable in the eyes of the law, and deemed protected classes. In the recent past, however, this definition has been expanded in a series of court judgements to include characteristics that are deemed difficult or too significant to ask someone to change. For instance, in June 2020, the U.S. Supreme Court issued a landmark judgement that found that discriminating against a person on the basis of their sexual orientation or gender identity in employment violates the protection for discrimination under sex in Title VII of the Civil Rights Act of 1964 [Bostock v. Clayton Cty., 140 S. Ct. 1731 (2020)].
@@
We argue that residential zip codes can also be considered an immutable characteristic for a section of the population, thereby deserving protection against discrimination on the basis of zip codes. Common sense dictates that it is not easy or straightforward for an individual or family to move to a lower-poverty or less segregated neighborhood in order to access better economic or educational opportunities. Even when disadvantaged families can afford to live in higher-income, more integrated neighborhoods with the support of housing vouchers they still cannot move due to source of income discrimination by landlords (Mazzara & Knudsen, 2019). Research has also shown that such discrimination overwhelmingly affects minorities, perpetuating systemic racism “and denies renters of color, especially Black renters, equal access to housing opportunities.” (Coa, 2023) Therefore, for low socioeconomic status families, where they live can be considered an immutable characteristic – they simply do not have the ability to move even if they wish or can afford to. Furthermore, they are forced to remain in their low-income neighborhoods which are likely to be correlated with race due to historical segregation, suffering the negative consequences associated with that in addition to the discrimination due to their source of income.
@@
While the fate of House Bill 2203 remains to be seen, the tangible impact on society of location’s correlation with protected attributes cannot be denied. Regardless of legal protections afforded to location bias, the fair-AI community has an ethical responsibility to strive to address it given the extensive impact our systems have on society. In this goal, computer scientists must contend with challenges specific to spatial data in addition to bias, as we shall see next.

### v2

# arXiv:2403.14040v2[cs.CY] 26 Feb 2026

@@ Abstract
Despite location being increasingly used in decision-making systems deployed in sensitive domains such as mortgages and insurance, little attention has been paid to the unfairness that may seep in due to the correlation of location with characteristics considered protected under anti-discrimination law, such as race or national origin. This position paper argues for the urgent need to consider fairness with respect to location, termed spatial fairness. It outlines the harms perpetuated through location’s correlation with protected characteristics, which may be particularly consequential due to its treatment as a neutral or purely technical attribute, abstracted from its historical, political, and socioeconomic context. This interdisciplinary work connects knowledge from fields such as public policy, economic development, and geography to highlight how existing fair-AI research falls short in addressing spatial biases, and fails to consider challenges unique to spatial data. Furthermore, we identify limitations in the small body of prior work on spatial fairness work, and propose guidelines to inform future research aimed at mitigating spatial biases in data-driven decision-making systems.
fairness machine learning, fair-AI, machine learning
@@ 1 Introduction
The adoption of data-driven decision-making to learn patterns from enormous amounts of historical data has surged in the last decade in socially impactful domains such as finance, housing, and public services, reducing human effort while increasing efficiency. However, as the fairness in AI (fair-AI) literature notes (Mehrabi et al., 2021; Pessach and Shmueli, 2022), such systems can replicate–or even amplify–undesirable patterns in historical data, such as racial bias, in ways that systematically disadvantage marginalized groups.
@@
This observation has led to research on developing fairer AI systems, centered on protecting against discrimination with respect to personal characteristics or traits protected by anti-discrimination law, commonly referred to in the literature as (legally) protected characteristics or attributes. However, even when this legislation is adhered to, discriminatory effects may persist because legally protected attributes can have proxies in the form of non-protected characteristics that are statistically or structurally entangled with them – such as location. Extensive research in public policy, economic development, and sociology has documented the extensive discrimination in the form of spatial discrimination that still adversely affects society today. For example, research has documented discrimination against racial minorities in the mortgage industry (Pager and Shepherd, 2008; Ross et al., 2008; Massey et al., 2016), despite the Equal Credit Opportunity Act prohibiting creditors from discriminating against applicants on the basis of race (Mehrabi et al., 2021). This is likely driven by the presence of extensive correlation of race with zip codes in the U.S. due to prejudiced practices such as redlining [^1] Redlining is the discriminatory practice of systematically denying mortgages, insurance, and other financial services to residents of specific neighborhoods based on their race or ethnicity(Woods, 2012; Rothstein, 2017). prevalent in the past (Schill and Wachter, 1994; Woods, 2012) rather than overt prejudice.
@@
The persistence of historical biases that disproportionately disadvantage racial and ethnic minorities (Rothstein, 2015; Schill and Wachter, 1994; Rothstein, 2017) in the outcomes of decision-making systems—despite legal protections for certain attributes—is alarming and warrants greater attention. While many attributes may be correlated with a legally protected attribute, the use of location in decision-making systems is especially consequential because it is often jointly correlated with multiple legally protected characteristics, such as race or national origin (Marshall, 1974; Willborn, 1984; Romei and Ruggieri, 2014; Becker, 2010). Table 1 lists protected characteristics inferable from location data.
Several approaches have been proposed to address the lack of consideration of legally protected attributes and their proxies in decision-making systems. Simply not considering protected attributes in decision-making, an approach known as ‘fairness through unawareness’ (Grgic-Hlaca et al., 2016; Kusner et al., 2017) has been shown to be insufficient in practice because historical data often encode discriminatory patterns through correlated features and downstream outputs (Lee and Floridi, 2021). Other fair-AI work has argued to “repair” or fix attributes that may act as proxies for legally protected characteristics (He et al., 2020; Kamiran et al., 2010; Berk et al., 2017; Zafar et al., 2017). However, subsequent research has shown this is ineffective in practice (Corbett-Davies and Goel, 2018; Lee and Floridi, 2021). Other work argues for dropping both proxies and protected characteristics from the data altogether (Datta et al., 2017), but this is not ideal as a proxy attribute can carry other information relevant to legitimate decision-making objectives, such as crime rates associated with zip codes in mortgage underwriting. These limitations of existing approaches motivate the need for fairness definitions and methods tailored to address spatial biases. Future work will need to overcome challenges unique to spatial data to address fairness concerns and extend the generalizability of existing fair-AI methods.
@@
This work advocates for a dedicated focus on spatial fairness within fair-AI. We use the interplay of legislation, historical discrimination, and AI systems within the U.S. to motivate greater attention towards bias due to location throughout the paper, however, the issue of discrimination due to location because of its correlation with other attributes is not limited to the U.S. alone (Galster and Friedrichs, 2015; Korsu and Wenglenski, 2010). By drawing on literature in the law, sociology, and economic development as well as computer science, we aim to ensure that future research by computer scientists in fair-AI does not fall into the trap of not being implementable in practice due to its noncompliance with legal or other standards (Xiang and Raji, 2019; Barocas and Selbst, 2016; Kim, 2022). We delineate challenges specific to location and draw upon relevant research in fields such as public policy and the law to establish guidelines for future research in spatial fairness. Our main contributions are:
@@
  • Motivate the importance of systematically addressing spatial fairness within the context of fair-AI, drawing upon scholarly literature from the law and public policy
  • Clearly outline challenges unique to spatial data
@@
  • Draw from fields of public policy, law, transport geography and economic development to delineate guidelines for future research on spatial fairness in fair-AI
Paper structure. Section 2 unpacks spatial bias, its societal impact, and the emerging legal landscape on zip-code discrimination. Sections 3 and 4 identify challenges unique to spatial data that limit fair-AI generalizability and limitations of current scholarship. We conclude with guidelines for future work in Section 5 and alternative views in Section 6.
@@
Race Federal
(Rothstein, 2015, 2017; Fulwood III, 2017)
National origin Federal
(Pandey and Caliskan, 2021; Ruther et al., 2018)
Source of income State
(Mazzara and Knudsen, 2019)
Protected in housing in 15 states (Policy et al., 2020)
Gender Federal
(Zhong et al., 2015; Riederer et al., 2016)
Table 1: Protected characteristics in housing (federal: Fair Housing Act; state: Policy Surveillance Program), credit (Equal Credit Opportunity Act, and employment (Equal Employment Opportunity Commission. (Mehrabi et al., 2021; Chen et al., 2019; Policy et al., 2020)
@@ 2 Context and Motivation: Location bias
We discuss spatial segregation in the U.S. and its lasting effects, and outline the case for legal protections against location-based discrimination, especially zipcodes.
@@ 2.1 Spatial unfairness in the real world
Although the fair-AI literature has yet to explore bias due to location in depth, it has been extensively covered in fields such as public policy, urban planning, and the law (Schill and Wachter, 1994; Rothstein, 2017; Baum et al., 2015). For example, neighborhoods in the U.S. have been historically correlated with race, a legally protected characteristic in many domains (e.g. housing, employment, credit), in complex ways due to prejudiced practices such as redlining. Despite being prohibited by the Fair Housing Act in 1968 [42 U.S.C. §§ 3601 - 3619], effects of redlining continue to linger (Angwin et al., 2017; Mehrabi et al., 2021). Zip codes in many parts of the country still correlate with race, potentially acting as a proxy for race in data-driven decision-making scenarios where zip codes are considered as a feature, such as in mortgages and insurance. This is well documented. For example, the non-profit ProPublica showed that drivers who live in minority neighborhoods are charged higher insurance premiums than drivers of same risk levels who reside in predominantly white neighborhoods (Angwin et al., 2017). Research has also found prices for ride-hailing trips in the city of Chicago to or from neighborhoods with larger non-white populations or higher poverty levels were significantly likely to be charged higher prices (Pandey and Caliskan, 2021; Saxena et al., 2023). Spatial bias has also been observed in commercial logistics algorithms. A Bloomberg study found Amazon Prime’s Same-Day Delivery service disproportionately excluded minority neighborhoods, and these spatial disparities could not be explained by income differences alone (Ingold and Soper, 2016).
@@
Neighborhoods (and their zip codes) may also be correlated with immigrants, i.e., regions where immigrants are more likely to live, which can act as a proxy for national origin, another protected characteristic in various domains under U.S. law (Mehrabi et al., 2021). Moreover, neighborhoods may also be correlated with income (Mode et al., 2016). Beyond the ride-hailing disparities noted above, low-income neighborhood residents may also face other forms of income-related discrimination. For example, families receiving housing vouchers often face barriers in moving to lower-poverty and higher-income neighborhoods from landlords who discriminate against their source of income (Mazzara and Knudsen, 2019). As a result, many families receiving housing vouchers still end up in low-income, racially segregated neighborhoods despite being able to afford homes in better areas (Rothstein, 2015). Such families might suffer doubly if they are also then subjected to spatial discrimination due to their low-income, racially segregated neighborhood. While not federally prohibited, several states have laws to protect from discrimination based on source of income (Table 1).
@@
Despite racial segregation in housing being banned in the U.S. in 1968, spatial segregation by race has persisted, resulting in “racially homogenous public institutions that are geographically defined,” (Fulwood III, 2017). As a consequence, such neighborhoods have reduced access to economic, educational, and social opportunities (Chetty et al., 2016; Li et al., 2013), which further hinders future prospects for residents in those neighborhoods. This also manifests in house prices, with a study finding that the racial composition of a neighborhood was a stronger determinant of a home’s appraised value in 2015 than in 1980 (Howell and Korver-Glenn, 2021). Furthermore, research has shown that prices for houses in majority-white neighborhoods are significantly higher than for similar houses in neighborhoods that happen to be majority-black (Rothwell and Perry, 2022), worsening the wealth gap between whites and racial minorities, especially African Americans, since a home is typically a family’s most significant financial asset according to a report by Pew Research Center (Taylor et al., 2011). There has been no significant change over the past few decades in the manner in which majority-black neighborhoods continue to be encircled by spatial disadvantage (Sharkey, 2014). Ethnographic research has documented that even middle class African-American families with an annual income of at least \$100,000 tend to live in neighborhoods with the same disadvantages as the average white family with an annual income of \$30,000 or lower (Sharkey, 2014).
@@
While our analysis leverages U.S. anti-discrimination law to motivate this problem, spatial inequities are a global issue. Research economics and public policy has documented rising spatial inequities in countries across the world (Kanbur and Venables, 2005; Kochendörfer-Lucius and Pleskovic, 2009; Zhao and Tong, 2000; Feldman and Storper, 2018).
@@ 2.2 Legal support for spatial fairness
U.S. law recognizes two forms of discrimination against protected classes (Lee and Floridi, 2021; Baum et al., 2015): disparate treatment and disparate impact. In layman terms, disparate treatment refers to direct or intentional discrimination while disparate impact refers to indirect or unintentional discrimination (Barocas and Selbst, 2016; Primus, 2010). While disparate treatment focuses on the intent to discriminate, disparate impact focuses on outcomes that are discriminatory towards members of a protected class despite a seemingly neutral decision-making policy.
@@
The law offers protections against both forms of discrimination on the basis of legally protected characteristics in domains such as housing [42 U.S.C. § 3601 et seq], credit [15 U.S.C. 1691 et seq.], and employment [42 U.S.C. § 2000e]. Some states define additional protected attributes: marital status is also a protected class in the state of California in the domain of employment (Gelb and Frankfurt, 1982) although it is a protected characteristic only in the domain of credit at the federal level. Typically under U.S. anti-discrimination law, protected characteristics are those which are considered immutable, i.e., characteristics that were not chosen and cannot be changed over time (Hoffman, 2010), such as race, color, and sex. Race and color were among the first classes to be considered immutable in the eyes of the law, and deemed protected classes. In the recent past, however, this definition has been expanded in a series of court judgments to include characteristics that are deemed difficult or too significant to ask someone to change. For instance, the U.S. Supreme Court issued a landmark judgment in June 2020 that found that discriminating against a person on the basis of their sexual orientation or gender identity in employment violates the protection for discrimination under sex in Title VII of the Civil Rights Act of 1964 [Bostock v. Clayton Cty., 140 S. Ct. 1731 (2020)].
@@
We argue that residential zip codes can also be considered an immutable characteristic for a section of the population, thereby deserving protection against discrimination on the basis of zip codes. Common sense dictates that it is not easy or straightforward for an individual or family to move to a lower-poverty or less segregated neighborhood in order to access better economic or educational opportunities. Even when disadvantaged families can afford to live in higher-income, more integrated neighborhoods with the support of housing vouchers they still cannot move due to source of income discrimination by landlords (Mazzara and Knudsen, 2019). Research has also shown that such discrimination overwhelmingly affects minorities, perpetuating systemic racism “and denies renters of color, especially Black renters, equal access to housing opportunities.” ((NLIHC), 2023) Therefore, for low socioeconomic status families, where they live can be considered an immutable characteristic – they simply do not have the ability to move even if they wish or can afford to. Furthermore, they are forced to remain in their low-income neighborhoods which are likely to be correlated with race due to historical segregation, suffering the negative consequences associated with that in addition to the discrimination due to their source of income.
@@
The tangible impact on society of location’s correlation with protected attributes is undeniable. Regardless of legal protections given to location, the AI community has an ethical responsibility to strive to address it given the extensive societal impacts of our systems. Doing so requires confronting challenges specific to spatial data, as we shall see next.